OpenStudy (anonymous):

could anyone explain why limit of [(1/t)-1 when t approaches 0+], is infinity? I don't really get the meaning of t approaching a+/a-. I understand it means approaching from left/right, but how's that related with infinity? Thanks a bunch!

OpenStudy (anonymous):

since infinity is an abstract concept, we often use approaching infinity to keep the integrity of a mathematical equation. For example if i said x+3y^2 = infinity, then the equation has lost its meaning. What do you assign to x and y to get infinity? similarly, for t = 0, 1/t = 1/0 which is infinity. So, to describe this, we say 1/t approaches infinity as t approaches 0. t can approach 0 from the right side, 0+ or the left side, 0-

OpenStudy (anonymous):

if it approaches from right side of the graph, you get + infinity, if it approaches from left side of graph, you get negative infinity. Infinity can stretch out in both directions.

OpenStudy (anonymous):

wait a sec, isn't 1/t just undefined when t=0? Why would it be infinite?

OpenStudy (anonymous):

how do you define infinity?

OpenStudy (anonymous):

if there is an axis infinity will go far far away?